OpenStudy (anonymous):

Solve: x2 - 20x + 36 = 0 a: {6, -6} b: {2, 18} c: {-4, -9} d: no solution

OpenStudy (welshfella):

you need 2 numbers a and b where a * b = +36 and a + b = -20

OpenStudy (welshfella):

these numbers go into the binomial (x + a)(x + b) = 0 then solve for x

OpenStudy (welshfella):

this equation will factor

OpenStudy (welshfella):

think of the factors of +36 4 * 9 -4*-9 2 * 18 -2 * -18 3 * 12 -3*-12 are some of them

OpenStudy (welshfella):

- which of those give you -10 when added?

OpenStudy (welshfella):

* -20

OpenStudy (anonymous):

You have been given points in your options, you can put them in your original equation and check which one is satisfying that equation..

OpenStudy (anonymous):

\[Given\ eq: \ x^2-20x+36\\ x^2-(18+2)x+36\ [ \because 36 \times1=36=18\times 2 |18+2=20]\\x^2-18x-2x+36\\x(x-18)-2(x-18)\\(x-18)(x-2)\\ \therefore\ original\ eq\ reduces\ as \ (x-18)(x-2)=0\\ \implies \ either\ x-18=0\ or\ x-2=0 \\ i.e.\ x=18 \ or\ x=2\\hence\ (b) is \ correct \ option.\]